# Vision Model Pre-training on Interleaved Image-Text Data via Latent Compression Learning

### v1

@@ Abstract
Recently, vision model pre-training has evolved from relying on manually annotated datasets to leveraging large-scale, web-crawled image-text data. Despite these advances, there is no pre-training method that effectively exploits the interleaved image-text data, which is very prevalent on the Internet. Inspired by the recent success of compression learning in natural language processing, we propose a novel vision model pre-training method called Latent Compression Learning (LCL) for interleaved image-text data. This method performs latent compression learning by maximizing the mutual information between the inputs and outputs of a causal attention model. The training objective can be decomposed into two basic tasks: 1) contrastive learning between visual representation and preceding context, and 2) generating subsequent text based on visual representation. Our experiments demonstrate that our method not only matches the performance of CLIP on paired pre-training datasets (e.g., LAION), but can also leverage interleaved pre-training data (e.g., MMC4) to learn robust visual representation from scratch, showcasing the potential of vision model pre-training with interleaved image-text data. Code is released at https://github.com/OpenGVLab/LCL.
@@ 1 Introduction
Figure 1: Comparison of different training frameworks. (a) shows the contrastive learning framework from CLIP [55], while (b) presents our novel LCL pre-training framework. Compared with CLIP, our LCL can effectively train vision encoders on interleaved image-text data. In these two frameworks, the text encoder or the language model that provides supervision can be optionally discarded during the transfer stage. (c) illustrates the multi-modal incremental training process, which can also be trained on interleaved image-text data. However, its primarily objective is to align the pre-trained vision encoder and the language model. Therefore, the pre-trained vision encoder in this framework can be selectively frozen or fine-tuned.
@@ 2 Related Work
Vision-centric Pre-training Methods. Supervised Pre-training on large-scale annotated datasets [24, 62, 48, 82] has remained the mainstream method for a long time, and has been favored by various vision tasks  [9, 24, 79, 7], demonstrating strong performance. Self-Supervised Pre-training has gained significant popularity due to its advantage of utilizing unlabeled data. BEiT  [6] follows the methodology of BERT [19] by randomly masking image tokens and reconstructing them as targets. MAE  [27] and SimMIM  [80] directly use masked pixels as reconstruction targets, making the pre-training process more straightforward and efficient. Weakly-Supervised Pre-training leverages image-hashtag  [50, 76, 63] and image-text datasets  [71, 61, 8, 57], which rely on noisy text supervision from the internet. For image-hashtag datasets, related works  [50, 63] have shown comparatively good performance across various transfer learning settings. In the case of image-text datasets, early efforts  [3, 42, 49, 64, 66, 67, 70, 13, 39] focused on learning general visual-linguistic representation. Recently, exemplified by CLIP  [55], methods  [55, 32] have been developed that involve pre-training through aligned visual-linguistic representation, achieving outstanding results in image classification tasks. Other works like M3I Pre-training  [65] propose a unified framework that integrates multiple pre-training strategies with data from various modalities and sources. Currently, weakly supervised pre-training from web-scale text supervision has become the core of multi-modal understanding, but existing methods have not yet to leverage the most widespread interleaved image-text data for training visual representation from scratch.
@@ 3 Method
Figure 2: Overview of our proposed Latent Compression Learning for vision model pre-training. Image latent representation is extracted via a vision encoder and subsequently has been input into a language model alongside textual embedding. Two complementary losses are utilized to learn robust visual representation from scratch on interleaved image-text data: a contrastive loss ensures consistency between the visual latent representation and its preceding context, while an auto-regressive loss enhances the predictability of visual representation for subsequent text.
@@ 3.1 Latent Compression Learning
where the $k$-th element of the output $y_{k}=g_{\phi}\circ f_{\theta}(x)_{k}$ predicts the next input latent $z_{k}$, $f_{\theta}$ is identical for text tokens for simplicity of annotation. When the compression only applied on text tokens, it degenerates to auto-regressive language modeling on interleaved image-text data used by previous methods (e.g., Kosmos [29] and Flamingo [2]).
@@
Here, we show that maximizing the mutual information between the output $y$ and the input latent $z$ of the causal attention model $g_{\phi}$ is equivalent to compressing $z$ by minimizing $H(p,q)$ in Eq. (2) meanwhile maximizing the entropy of each element $z_{k}$ in $z$:
@@ 3.2 Training Loss
In this sub-section, we demonstrate how Eq. (3.1) is decomposed into training tasks and losses. Firstly, $I(y;z)$ can be decomposed as a cross-entropy term and an entropy term in the following two symmetric ways (see Appendix for detailed derivation):
@@
Relation to Previous Pre-training Tasks. In our proposed pre-training framework, the contrastive task and generation task align the image representation with preceding context and subsequent context, respectively. Hence, combining these two pre-training tasks can fully leverage the semantic information contained in interleaved image-text data to supervise the learning of image representation. For previous pre-training tasks, 1) Contrastive Language-Image Pre-training (CLIP) [55] has a similar objective of maximizing the mutual information between corresponding image and text [65], but it can only be applied to paired image-text data. 2) Auto-regressive Text Generation task only leverages the semantic information in subsequent context to supervise the learning of image representation, while the preceding context is missing, and representation collapse cannot be avoided. Moreover, interleaved image-text data usually has information redundancy, i.e., the image and its corresponding text may contain similar information. So models may rely on information from the text rather than the image for prediction. This is particularly true when the vision encoder is trained from scratch, resulting in the image representation that is never focused or optimized. Our experiments in Sec. 4.2 confirms this analyse.
@@ 3.3 Architecture
The overview of model architecture when adopting our LCL is shown in Fig. 2. In the interleaved image-text input sequence, we introduce special tokens <BoI> and <EoI> to serve as special tokens for the beginning and the ending of the visual embeddings from an image, respectively. We adopt a Vision Transformer (ViT) [20] as the vision encoder, which inputs a sequence of image patches and outputs a sequence of latent representation.
In the contrastive learning task, calculating the loss for each image token is extremely computationally expensive. To alleviate this problem, we consider utilizing one global representation per image instead of all latent representation for contrastive learning. Specifically, the global representation $v_{i}$ of each image is extracted from its latent representation $z_{i:i+M}$ through an attention pooling, which is a multi-head attention layer with a learnable query token. The output of the causal transformer model at <BoI> token, just before the image, is processed by a LayerNorm layer and a linear projection into $t_{i}$. The contrastive loss in Eq. (6) is calculated between $v_{i}$ and $t_{i}$. The text generation loss is the standard cross-entropy loss over the output logits of the language model defined in Eq. (7).
@@ 4.1 Experiment Settings
Pre-train Data. The datasets utilized in our pre-training encompass the image-text pair dataset LAION-400M [57], as well as the image-text interleaved datasets MMC4 [88] and OBELICS [36]. We also re-organize two datasets for fair comparison with CLIP. 1) LAION-Random as a interleaved dataset. Images from LAION-400M are randomly placed before or after their paired caption to form image-text sequences. 2) MMC4-Pair as a paired dataset. For images in MMC4, we select matching text from the bipartite graph matching results derived from CLIP similarity to generate pseudo image-text paired data.
Implementation Details. We adopt the same image transform in OpenCLIP [15] for pre-training and set the image size as $224\times 224$ for all experiments. We employ ViT [20] as our vision encoder. ViT-L/14 is used for the main results, and ViT-B/16 is used for ablation studies. The language model follows the same architecture as OPT-125M [84] but is randomly initialized. By default, the contrastive loss balancing weight is set at $\lambda=0.1$. AdamW optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.95$ and a weight decay of 0.1 are used. We employ a cosine learning rate schedule with linear warmup and set the peak learning rate at 5e-4 for the LAION data and 3e-4 for others. The model for the main results is trained for 200k iterations with 64k images per batch on average. In the ablation study, the models are trained for 25k iterations with 8k images per batch.
@@
* Pre-training method Pre-training data IN-1k COCO ret. Flickr30k ret. COCO cap. NoCaps cap.
@@
Con. (CLIP) MMC4-Pair 74.8 46.4 32.5 76.2 60.0 23.9 82.9 29.6 78.4
Con. + Cap. (CoCa) MMC4-Pair 75.4 48.6 34.3 76.5 61.9 23.7 84.8 30.0 80.5
Con. + Cap. + Mat. (BLIP2) MMC4-Pair 74.5 46.5 31.3 74.9 57.8 23.7 82.9 29.4 78.1
Mask. (BEiT3) MMC4 73.3 45.1 30.6 73.2 57.1 23.3 81.4 29.5 76.7
Gen. (Flamingo) MMC4 24.0 10.6 5.6 17.7 10.8 8.7 18.1 15.0 18.5
Gen. + Reg. (Emu) MMC4 5.7 2.3 1.4 4.8 2.7 0.3 4.4 0.6 4.6
@@
Con. (CLIP) LAION-400M 75.0 47.2 34.2 76.5 59.8 24.1 84.1 30.0 78.8
Con. + Cap. (CoCa) LAION-400M 75.2 48.6 34.8 76.9 61.4 24.6 88.2 30.4 82.9
Con. + Cap. + Mat. (BLIP2) LAION-400M 74.0 47.4 31.5 75.9 57.9 23.6 85.0 30.0 77.8
@@
4.2 Ablation Study
We divide existing vision model pre-training methods into the following representative pre-training tasks: (1) Image-text contrastive (Con.) [55, 15, 32], (2) Image-text contrastive + image captioning (Con. + Cap.) [81], (3) Image-text contrastive + image captioning + image-text matching (Con. + Cap. + Mat.) [40, 41, 14], (4) Auto-regressive text generation (Gen.) [2, 29, 44], (5) Auto-regressive text generation + image regression (Gen. + Reg.) [69, 68, 72], (6) Mask data modeling (Mask.) [78].
Comparing Vision Model Pre-training Methods on Interleaved Image-Text Data We conduct experiments on the MMC4 dataset [88] to demonstrate the effectiveness of our LCL on interleaved image-text data. For pre-training methods that only support paired data, MMC4-Pair is used.
@@ 4.2 Ablation Study
Comparing Vision Model Pre-training Methods on Paired Image-Text Data. We conduct experiments on the LAION-400M dataset to show that LCL pre-training also performs well on paired image-text data without specific modification. Tab 2 shows that our method is comparable to paired pre-training methods on various tasks, indicating that all information in paired data is fully exploited.
@@
Table 4: Ablations of the training loss and loss balancing weight in LCL .
@@
4.3 Comparison with OpenCLIP
Table 5: Comparison with OpenCLIP on classification, retrieval and caption tasks with ViT/L-14.
acc-1 TR@1 IR@1 TR@1 IR@1 B@4 C B@4 C
frozen transfer
OpenAI CLIP WIT-400M 32 83.7 61.7 48.2 89.0 75.8 32.1 116.0 35.5 108.9
OpenCLIP LAION-400M 32 82.1 59.5 46.0 86.9 74.2 31.0 111.5 34.8 106.0
LCL (Ours) LAION-400M 32 82.2 59.6 46.2 86.7 74.0 31.4 112.3 35.0 106.7
full transfer
OpenAI CLIP WIT-400M 32 87.4 62.1 49.6 90.3 77.9 39.5 132.7 41.4 116.9
OpenCLIP LAION-400M 32 86.2 61.7 47.5 87.7 76.3 38.6 128.9 39.9 112.7
LCL (Ours) LAION-400M 32 86.1 61.9 47.5 87.6 76.1 39.1 129.7 40.2 113.5
Table 6: Comparison with OpenCLIP on multi-modal benchmarks with ViT/L-14.
Model Pre-training data Pre-training epoch VQAv2 GQA VisWiz SQA POPE MME MMB SEED
frozen transfer
OpenAI CLIP WIT-400M 32 77.1 61.7 44.4 71.1 84.6 1486.9 65.1 64.6
OpenCLIP LAION-400M 32 68.7 57.0 39.5 69.0 81.8 1266.2 54.2 55.8
full transfer
OpenAI CLIP WIT-400M 32 79.0 62.8 46.8 71.8 85.7 1576.8 68.9 67.9
OpenCLIP LAION-400M 32 71.5 58.6 42.2 70.4 82.5 1345.4 58.5 59.5
We use our proposed Latent Compression Learning (LCL ) to pre-train a ViT-L/14 with mixed data from the LAION-400M and MMC4. We compare it to the ViT-L/14 pre-trained by OpenCLIP [15] using the public LAION-400M dataset. For a fair comparison, the total number of images seen during pre-training is 13B. Tab. 5 and Tab. 6 show the results. When both use LAION-400M as pre-training data, as demonstrated by previous experimental conclusions, LCL achieves similar performance to CLIP. When combined with MMC4, our method outperforms CLIP, especially in caption and multi-modal dialogue tasks.

### v2

@@ Abstract
Recently, vision model pre-training has evolved from relying on manually annotated datasets to leveraging large-scale, web-crawled image-text data. Despite these advances, there is no pre-training method that effectively exploits the interleaved image-text data, which is very prevalent on the Internet. Inspired by the recent success of compression learning in natural language processing, we propose a novel vision model pre-training method called Latent Compression Learning (LCL) for interleaved image-text data. This method performs latent compression learning by maximizing the mutual information between the inputs and outputs of a causal attention model. The training objective can be decomposed into two basic tasks: 1) contrastive learning between visual representation and preceding context, and 2) generating subsequent text based on visual representation. Our experiments demonstrate that our method not only matches the performance of CLIP on paired pre-training datasets (e.g., LAION), but can also leverage interleaved pre-training data (e.g., MMC4) to learn robust visual representations from scratch, showcasing the potential of vision model pre-training with interleaved image-text data. Code is released at https://github.com/OpenGVLab/LCL.
@@ 1 Introduction
Figure 1: Comparison of different training frameworks. (a) Contrastive learning framework from CLIP [55] pre-trains vision encoders from scratch with image-text pairs, but it does not support interleaved data. (b) Our proposed LCL pre-training frameworkm can effectively pre-train vision encoders from scratch with interleaved image-text data. In these two frameworks, the text encoder or the language model that provides supervision can be optionally discarded during the transfer stage. (c) Multi-modal incremental training process uses interleaved image-text data to align the pre-trained vision encoder and the language model, but it cannot pre-train vision encoders from scratch.
@@ 2 Related Work
Vision-centric Pre-training Methods. Supervised Pre-training on large-scale annotated datasets [24, 62, 48, 82] has remained the mainstream method for a long time, and has been favored by various vision tasks  [9, 24, 79, 7] demonstrating strong performance. Self-Supervised Pre-training has gained significant popularity due to its advantage of utilizing unlabeled data. BEiT  [6] follows the methodology of BERT [19] by randomly masking image tokens and reconstructing them as targets. MAE  [27] and SimMIM  [80] directly use masked pixels as reconstruction targets, making the pre-training process more straightforward and efficient. Weakly-Supervised Pre-training leverages image-hashtag  [50, 76, 63] and image-text datasets  [71, 61, 8, 57], which rely on noisy text supervision from the internet. For image-hashtag datasets, related works  [50, 63] have shown comparatively good performance across various transfer learning settings. In the case of image-text datasets, early efforts  [3, 42, 49, 64, 66, 67, 70, 13, 39] focused on learning general visual-linguistic representation. Recently, exemplified by CLIP  [55], methods  [55, 32] have been developed that involve pre-training through aligned visual-linguistic representation, achieving outstanding results in image classification tasks. Other works like M3I Pre-training  [65] propose a unified framework that integrates multiple pre-training strategies with data from various modalities and sources. Currently, weakly supervised pre-training from web-scale text supervision has become the core of multi-modal understanding, but existing methods have not yet to leverage the most widespread interleaved image-text data for training visual representation from scratch.
@@ 3.1 Latent Compression Learning
where the $k$-th element of the output $y_{k}=g_{\phi}\circ f_{\theta}(x)_{k}$ predicts the next input latent $z_{k}$, $f_{\theta}$ is identity for text tokens for simplicity of annotation. When the compression only applied on text tokens, it degenerates to auto-regressive language modeling on interleaved image-text data used by previous methods (e.g., Kosmos and Flamingo).
@@
Prior work [65] have shown the relationship between cross-entropy and mutual information in other pre-training tasks. Here, we derive this relationship in the latent compression task: maximizing the mutual information between the output $y$ and the input latent $z$ of the causal attention model $g_{\phi}$ is equivalent to compressing $z$ by minimizing $H(p,q)$ in Eq. (2) meanwhile maximizing the entropy of each element $z_{k}$ in $z$:
@@ 3.2 Training Loss
In this sub-section, we demonstrate how Eq. (3.1) is decomposed into training tasks and losses. Firstly, $I(y;z)$ can be decomposed as a cross-entropy term and an entropy term in the following two symmetric ways (see Appendix B for detailed derivation):
@@
Relation to Previous Pre-training Tasks. In our proposed pre-training framework, the contrastive task and generation task align the image representation with preceding context and subsequent context, respectively. Hence, combining these two pre-training tasks can fully leverage the semantic information contained in interleaved image-text data to supervise the learning of image representation. For previous pre-training tasks, 1) Contrastive Language-Image Pre-training (CLIP) [55] has a similar objective of maximizing the mutual information between corresponding image and text [65], but it can only be applied to paired image-text data. 2) Contrastive Captioner (CoCa) [81] combines a captioning (text generation) loss with the CLIP loss, but it cannot be applied to interleaved data, either. CLIP loss requires image-text pairs, while the captioner can only be conditioned on a single image can rather than flexible interleaved image-text contents. 3) Auto-regressive Text Generation task only leverages the semantic information in subsequent context to supervise the learning of image representation, while the preceding context is missing, and representation collapse cannot be avoided. Moreover, interleaved image-text data usually has information redundancy, i.e., the image and its corresponding text may contain similar information. So models may rely on information from the text rather than the image for prediction. This is particularly true when the vision encoder is trained from scratch, resulting that the image representation are never focused and optimized. Our experiments in Sec. 4.2 confirms this analyze.
Figure 2: Overview of our proposed Latent Compression Learning for vision model pre-training. Image latent representation is extracted via a vision encoder and subsequently input into a language model alongside textual embedding. Two complementary losses are utilized to learn robust visual representation from scratch on interleaved image-text data: a contrastive loss ensures consistency between the visual latent representation and its preceding context, while an auto-regressive loss enhances the predictability of visual representation for subsequent text.
@@ 3.3 Architecture
The overview of model architecture when adopting our LCL is shown in Fig. 2. The interleaved image-text input sequence may contain multiple images. We adopt a Vision Transformer (ViT) [20] as the vision encoder, which encodes each image into a sequence of visual embeddings as its latent representation. The visual embeddings of each image are inserted at corresponding positions in the interleaved sequence, and we introduce special tokens <BoI> and <EoI> to indicate the beginning and the ending positions, respectively. The combined visual and text embeddings are fed into a causal language model. The text generation loss is the standard cross-entropy loss over the output logits of the language model defined in Eq. (7). In the contrastive learning task, calculating the loss for each image token is extremely computationally expensive. To alleviate this problem, we consider utilizing one global representation per image instead of all latent representation for contrastive learning. Specifically, the global representation $v_{i}$ of each image is extracted from its latent representation $z_{i:i+M}$ through an attention pooling, which is a multi-head attention layer with a learnable query token. The output of the causal transformer model at <BoI> token, just before the image, is processed by a LayerNorm layer and a linear projection into $t_{i}$. The contrastive loss in Eq. (6) is calculated between $v_{i}$ and $t_{i}$.
@@ 4.1 Experiment Settings
Pre-train Data. The datasets utilized in our pre-training encompass the image-text pair dataset LAION-400M [57], as well as the image-text interleaved datasets MMC4 [88] and OBELICS [36].
We also re-organize two datasets for fair comparison with CLIP. 1) LAION-Random as a interleaved dataset. Images from LAION-400M are randomly placed before or after their paired caption to form image-text sequences. 2) MMC4-Pair as a paired dataset. For images in MMC4, we select matching text from the bipartite graph matching results derived from CLIP similarity to generate pseudo image-text paired data.
Implementation Details. We adopt the same image transform in OpenCLIP [15] for pre-training and set the image size as $224\times 224$ for all experiments. We employ ViT [20] as our vision encoder. ViT-L/14 is used for the main results, and ViT-B/16 is used for ablation studies. The language model follows the same architecture as OPT-125M [84] but is randomly initialized. By default, the contrastive loss balancing weight is set at $\lambda=0.1$. AdamW optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.95$ and a weight decay of 0.1 are used. We employ a cosine learning rate schedule with linear warmup and set the peak learning rate at 5e-4 for the LAION data and 3e-4 for others. The model for the main results is trained for 200k iterations with 64k images per batch on average. In the ablation study, the models are trained for 250k iterations with 8k images per batch.
@@
* Pre-training method (task) Pre-training data IN-1k COCO ret. Flickr30k ret. COCO cap. NoCaps cap.
@@
CLIP (Con.) MMC4-Pair 74.8 46.4 32.5 76.2 60.0 23.9 82.9 29.6 78.4
CoCa (Con. + Cap.) MMC4-Pair 75.4 48.6 34.3 76.5 61.9 23.7 84.8 30.0 80.5
BLIP2 (Con. + Cap. + Mat.) MMC4-Pair 74.5 46.5 31.3 74.9 57.8 23.7 82.9 29.4 78.1
BEiT3 (Mask.) MMC4 73.3 45.1 30.6 73.2 57.1 23.3 81.4 29.5 76.7
Flamingo (Gen.) MMC4 24.0 10.6 5.6 17.7 10.8 8.7 18.1 15.0 18.5
Emu (Gen. + Reg.) MMC4 5.7 2.3 1.4 4.8 2.7 0.3 4.4 0.6 4.6
@@
* Pre-training method (task) Pre-training data IN-1k COCO ret. Flickr30k ret. COCO cap. NoCaps cap.
@@
CLIP (Con.) LAION-400M 75.0 47.2 34.2 76.5 59.8 24.1 84.1 30.0 78.8
CoCa (Con. + Cap.) LAION-400M 75.2 48.6 34.8 76.9 61.4 24.6 88.2 30.4 82.9
BLIP2 (Con. + Cap. + Mat.) LAION-400M 74.0 47.4 31.5 75.9 57.9 23.6 85.0 30.0 77.8
@@
4.2 Comparison with Vision Pre-training Methods
In this section, we show the superiority of our LCL when using interleaved image-text data for vision pre-training. However, existing vision pre-training methods only support paired data, so we choose MMC4 dataset with a paired set (MMC4-Pair) available for those methods. In addition, to confirm that it is is not trivial to use interleaved data for vision pre-training, we also consider existing methods to use interleaved data. Those methods are originally proposed as MLLM training methods, but here we test them for vision pre-training. Besides, on paired image-text data, we also compare our LCL with existing vision pre-training methods and show they are comparable. See Appendix A.3 and A.2 for detailed pre-training and evaluation settings.
We aggregate all the pre-training methods divide them into the following tasks: (1) Image-text contrastive (Con.) [55, 15, 32], (2) Image-text contrastive + image captioning (Con. + Cap.) [81], (3) Image-text contrastive + image captioning + image-text matching (Con. + Cap. + Mat.) [40, 41, 14], (4) Auto-regressive text generation (Gen.) [2, 29, 44], (5) Auto-regressive text generation + image regression (Gen. + Reg.) [69, 68, 72], (6) Mask data modeling (Mask.) [78]. We select one representative method from each task. For fair comparison, we implement these methods on the same model and training data with available open-source codes or our reproduction.
Pre-training on Interleaved Image-Text Data. We conduct experiments on the MMC4 dataset [88] to demonstrate the effectiveness of our LCL on interleaved image-text data. For pre-training methods that only support paired data, MMC4-Pair is used.
@@ 4.2 Comparison with Vision Pre-training Methods
Pre-training on Paired Image-Text Data. We conduct experiments on the LAION-400M dataset to show that LCL pre-training also performs well on paired image-text data without specific modification. Tab 2 shows that our method is comparable to paired pre-training methods on various tasks, indicating that all information in paired data is fully exploited.
4.3 Comparison with Pre-trained Checkpoints
Table 3: Transfer evaluation results of pre-trained ViT/L-14 on classification, retrieval and captioning tasks.
acc-1 TR@1 IR@1 TR@1 IR@1 B@4 C B@4 C
frozen transfer
OpenAI CLIP WIT-400M 32 83.7 61.7 48.2 89.0 75.8 32.1 116.0 35.5 108.9
OpenCLIP LAION-400M 32 82.1 59.5 46.0 86.9 74.2 31.0 111.5 34.8 106.0
LCL (Ours) LAION-400M 32 82.2 59.6 46.2 86.7 74.0 31.4 112.3 35.0 106.7
full transfer
OpenAI CLIP WIT-400M 32 87.4 62.1 49.6 90.3 77.9 39.5 132.7 41.4 116.9
OpenCLIP LAION-400M 32 86.2 61.7 47.5 87.7 76.3 38.6 128.9 39.9 112.7
LCL (Ours) LAION-400M 32 86.1 61.9 47.5 87.6 76.1 39.1 129.7 40.2 113.5
Table 4: Transfer evaluation results of pre-trained ViT/L-14 on multi-modal benchmarks. The transfer adopts the downstream model and training pipeline of LLaVA-1.5 [46].
Model Pre-training data Pre-training epoch VQAv2 GQA VisWiz SQA POPE MME MMB $\text{SEED}_{\text{I}}$
frozen transfer
OpenAI CLIP WIT-400M 32 77.1 61.7 44.4 71.1 84.6 1486.9 65.1 64.6
OpenCLIP LAION-400M 32 68.7 57.0 39.5 69.0 81.8 1266.2 54.2 55.8
full transfer
OpenAI CLIP WIT-400M 32 79.0 62.8 46.8 71.8 85.7 1576.8 68.9 67.9
OpenCLIP LAION-400M 32 71.5 58.6 42.2 70.4 82.5 1345.4 58.5 59.5
To further confirm the effectiveness of our proposed Latent Compression Learning (LCL), we compare our pre-trained model with existing checkpoints of pre-trained vision encoders. We use LCL to pre-train a ViT-L/14 with mixed data from the LAION-400M and MMC4. We compare it to the ViT-L/14 pre-trained by OpenCLIP [15] using the public LAION-400M dataset, and the ViT-L/14 pre-trained by OpenAI CLIP [55] with private data is listed as reference. The total number of images seen during pre-training is 13B for all models. We evaluate the pre-trained vision encoders by transferring to downstream tasks, i.e., integrate the vision encoders into downstream task models and compare the fine-tuning results. More training details are in Appendix A.1 and evaluation details are in Appendix A.2.
Tab. 3 and Tab. 4 show the results of transfer evaluations. When both use LAION-400M as pre-training data, as with the previous experimental conclusions, LCL has similar performance to CLIP. When combined with MMC4, our method achieves better performance, especially on caption and multi-modal dialogue tasks.
There exist approaches achieving better results on some benchmarks. However, they either use larger vision encoders, more training data, or private data. We list those results as reference in Appendix C.
4.4 Ablation Study
@@
Table 6: Ablations of the training loss and loss balancing weight in LCL . Models are evaluated under frozen transfer setting.